# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v3

@@ I Introduction
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works [5, 6, 7, 8, 9] exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [10, 11] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[12], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g. transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table I). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
The contributions of this work can be summarized as follows:
  • We systemically investigate the adversarial robustness of state-of-the-art AIGI detectors, including standard trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
  • We propose a novel attack against AIGI detection by exploring the vulnerable frequency region in a Bayesian manner. A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
  • We conduct comprehensive experiments on 13 various AIGI detectors, including heterogeneous model architectures and various detection methods. Extensive results demonstrate that our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
@@ II Related Work
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs [1], VAEs [13], and Diffusion models [2], has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods [14, 15, 16], to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches [3, 17] have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods [10, 11] focus on analyzing forgery traces in the frequency domain and learning relevant forgery features [18] to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features [19, 20] in the spatial domain. AIGI fingerprint representations [21] are acquired through the learning of noise patterns. Notably, studies such as [22, 4] have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies [23] has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD [24] explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error [25, 26], adversarial teacher-student discrepancy-aware framework [27] or diffusion noise[28]. The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including image classification[29, 30], multimedia communication [31] and human activity recognition [32]. Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. [6] employ gradient sign-based methods [33, 34] to evaluate the adversarial robustness of video Deepfake detectors. Li et al. [35] generate adversarial fake images along the latent face manifold. Carlini et al. [5] and Neekhara et al. [7] investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. [9] propose to generate adversarial perturbations in the frequency domain. Hou et al. [8] narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations. FakePolisher [36] and TraceEvader [37] improve the untraceability of deepfakes by removing artificial fingerprints. Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
@@ III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 21, 4]. While these subtle clues are invisible in the spatial domain, a series of works [10, 11] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [39] to visualize the sensitive components of real and fake images across different models:
$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
@@
$\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
@@ III-D Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3, 21]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
@@
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
@@ IV-A Experimental Settings
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43]. GenImage [12] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [12], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46]. The Deepfake dataset consists of 50k real face images from FFHQ [47] and 50k generated face images generated by StyleGAN2 [47]. After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 13 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [48], EfficientNet [49] and DenseNet [50], and ViT-based detectors Vision Transformer(ViT) [51] and Swin-Transformer(Swin-ViT) [52]. For evaluating on various detection methods, we use frequency-based detectors DCTA [11] and Spec [18], gradient-based detectors LGrad [23], CLIP-based detectors UnivFD [24], and diffusion-based detectors DNF [28]. In addition, LNP [22] extracts the noise pattern of images and GramNet [53] learns the global texture representation. We also consider them as victim models.
@@
Compared Methods: We adopt gradient-based methods, I-FGSM [33], PGD [38], MI [54] and frequency-based methods S2[39], SSAH [55] and ensemble-based methods ENS [54] and SVRE [56]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks Fakepolisher [36] and TraceEvader [37] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
@@
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3]. Specificaly, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3]. On the FFHQ (StyleGAN2) dataset we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [57]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [39]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
TABLE I: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
@@
LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
@@
GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
@@ IV-B Attack on Spatial-based and Frequency-based Detectors
We report the attack performance against spatial-based and frequency-based detectors in Table I. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.6%, while S2I, PGD, IFGSM and MIFGSM only have 93.1%, 93.1%, 66.2% and 66.2%. Although SSAH has a relatively high white-box success rate (97.7%), it fails to transfer the adversarial examples to black-box models. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 50.5%, surpassing the IFGSM, MIFGSM, PGD, S2I and SSAH by a big margin of 25.7%, 23.3%, 8.4%, 11.9% and 49.1%. Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
TABLE II: The attack success rate(%) compared with attacks agaisnt Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.

### v4

@@ I Introduction
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging. Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [5, 6], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts. This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints. Consequently, transferring adversarial patterns across such diverse representations is far less straightforward. Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [7, 8, 9, 10, 11]. These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces. In contrast, AIGI detectors are designed to identify any AI-generated content produced by a wide range of generative techniques, including GANs, diffusion models, and autoregressive models. The content extends beyond human faces to encompass non-human entities, landscapes, abstract art, and more. These differences make designing attacks for AIGI detection more challenging than Deepfake detection.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [12, 13] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[14], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g. transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
The contributions of this work can be summarized as follows: (1) We systematically assess the adversarial robustness of state-of-the-art AIGI detectors, revealing both real-world threats and gradient masking effects. (2) We propose a new attack against AIGI detection by exploring the vulnerable frequency region using a post-train Bayesian strategy, avoiding retraining the surrogate model. (3) Extensive experiments across 17 AIGI detectors show that our method achieves the highest attack success rates under white-box and black-box settings, outperforming existing baselines.
@@ II Related Work
AI-Generated Image Detection: The rapid progress of generative models, from early GANs [1] to recent diffusion models [2], has raised substantial risks of disinformation. While early work focused primarily on detecting fake human faces [15, 16, 17, 18], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection. AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images. Data-driven methods [3] achieve strong in-distribution performance but generalize poorly to unseen generators. To address this, frequency-domain approaches [13, 19] exploit forgery traces, while spatial-domain methods learn local forgery features [20]. Others leverage fingerprint representations from noise patterns [21, 4]. Parameter-efficient strategies based on frozen pre-trained models have also been explored [22, 23], along with approaches exploiting diffusion reconstruction error [24, 25], teacher–student discrepancies [26], or diffusion noise [27].
Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [28, 5], multimedia communication [29], and human activity recognition [30]. Gradient-based attack methods [5, 31] exploit gradients to mislead models. To improve adversarial transferability, frequency-domain approaches [32, 6] perturb spectral representations, while ensemble-based attacks [33, 34] aggregate gradients or perturbations from multiple models. Very recently, their impact on Deepfake detection has been explored. Gradient sign-based attacks [8, 5, 35], latent space manipulations [36], and black-box evaluations [7, 9] reveal vulnerabilities in existing detectors. Other works leverage frequency-domain perturbations [11], natural degradation noise [10], or fingerprint removal strategies such as FakePolisher [37] and TraceEvader [38]. Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [39]. In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
@@ III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 4]. While these subtle clues are invisible in the spatial domain, a series of works [13] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [32] to visualize the sensitive components of real and fake images across different models:
$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
@@
$\Gamma(\mathbf{x})=\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
@@ III-D Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
@@
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
@@ IV-A Experimental Settings
Datasets: We chose three generated image datasets created by a wide range of generative models. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43]. GenImage [14] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [14], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46]. The Deepfake dataset consists of 50k real face images from FFHQ [1] and 50k generated face images generated by StyleGAN2 [1]. After training, we collect only the correctly classified testing samples for attack in evaluation.
TABLE I: Datasets employed for training each detector.
Synthetic LSUN GenImage Synthetic FFHQ
CNNSpot
DenseNet
EfficientNet
MobileNet
Spec
DCTA
ViT
Swin
GramNet - -
LGrad - -
LNP - -
UnivFD - -
DNF - -
FreqNet - -
FreqMask - -
DIRE - -
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [47], EfficientNet [48] and DenseNet [49], and ViT-based detectors Vision Transformer(ViT) [50] and Swin-Transformer(Swin-ViT) [51]. For evaluating on various detection methods, we use frequency-based detectors DCTA [13], Spec [19], FreqNet [52], and FreqMask [53], gradient-based detectors LGrad [22], CLIP-based detectors UnivFD [23], and diffusion-based detectors DNF [27], DIRE [24] and AEROBLADE [25]. In addition, LNP [21] extracts the noise pattern of images and GramNet [54] learns the global texture representation. We also consider them as victim models. The datasets employed for training each detector are summarized in Table I. Note that AEROBLADE is a training-free method and not trained on the listed datasets.
@@
Compared Methods: We adopt gradient-based methods, I-FGSM [5], PGD [31], MI [33] and frequency-based methods S2[32], SSAH [6] and ensemble-based methods ENS [33] and SVRE [34]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [11], Fakepolisher [37] and TraceEvader [38] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
@@
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3]. Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3]. On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [55]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [32]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
TABLE II: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
@@
Synthetic LSUN (ProGAN)
FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
@@
GenImage(SD)
FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
@@ IV-B Attack on Spatial-based and Frequency-based Detectors
We report the attack performance against spatial-based and frequency-based detectors in Table II. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.6%, while S2I, PGD, IFGSM and MIFGSM only have 93.1%, 93.1%, 66.2% and 66.2%. Although SSAH has a relatively high white-box success rate (97.7%), it fails to transfer the adversarial examples to black-box models. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 50.5%, surpassing the IFGSM, MIFGSM, PGD, S2I and SSAH by a big margin of 25.7%, 23.3%, 8.4%, 11.9% and 49.1%. Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
TABLE III: The attack success rate(%) compared with attacks against Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.